Freshman Seminar 21n: Elliptic Curve Projects

William Stein

April 8, 2003

Today you will give initial presentations about your projects. Your assignment is to write a four paragraph email to me, with one paragraph about each of the other projects being done by the other students in the seminar. My initial impression is that Mauro is doing a historical survey of cubics, Jeff and Alex are studying the Birch and Swinnerton-Dyer conjecture, Jennifer is figuring out how Michael Stoll found a generator for the Mordell-Weil group of y^{2} = x^{3}+ 7823, and Jenna is writing about modularity of elliptic curves.

1 Schedule

April 15 Second presentation
April 22 Third presentation
April 29 Fourth presentation (fairly polished draft due)
May 6 Fifth presentation (get comments on your polished draft)
May 13 Give a polished 40 minute presentation

Note: I will be gone on April 8 and May 5–6 (Institute for Defense Analysis), but Grigor will lead class those two days. I will be easy to reach via email.